What is the Power in dipoters of diverging lens of focal length 0.4m,0.1m, 80mm


1
Expert's answer
2021-06-11T09:46:57-0400

Power of lens of focal length 0.4 m = "\dfrac{1}{F(in\space m)}=\dfrac{1}{-0.4}=-2.5\space D"

Power of lens of focal length 0.1 m = "\dfrac{1}{F(in\space m)}=\dfrac{1}{-0.1}=-10\space D"

Power of lens of focal length 80 mm(0.08 m) = "\dfrac{1}{F(in\space m)}=\dfrac{1}{-0.08}=-12.5\space D"